OpenStudy (anonymous):

I am having trouble with dividing scientific notation could anyone help me with this problem (8.56 · 10^8) / (3.71 · 10^3) I don't just want the answer i actually want to learn how to do it so I can figure out what I am doing wrong

OpenStudy (anonymous):

You can write the numerator as (8.56)(10^8) and the denomimantor as (3.71)(10^3) So you can divide the number and exponents Since its 10^8 / 10^3, you subtract the exponents, and you just divide the numbers normally You get ((8.56)(10^5))

OpenStudy (anonymous):

ok how did you get the 8.56 and where did the 3.71 go?

OpenStudy (anonymous):

my bad, i meant to write 8.56/ 3.71

OpenStudy (anonymous):

((8.56/3.71)(10^5))

OpenStudy (anonymous):

ok then to simplify this do I divide those because then I get a really long yucky decimal?

OpenStudy (anonymous):

Yeah, when you try to simplify, you get a long decimal number

OpenStudy (anonymous):

so that would be the simplified form to avoid the long decimal?

OpenStudy (mertsj):

\[\frac{8.56\times10^{8}}{3.71\times10^{3}}=\frac{8.56}{3.71}\times\frac{10^{8}}{10^{3}}\]

OpenStudy (mertsj):

\[2.31\times10^{5}\]

OpenStudy (anonymous):

so you are saying that you can just simplify the long decimal to 2.31 to simplify the whole problem?

OpenStudy (anonymous):

If you write 2.31 its only ABOUT 8.56/3.71, not exactly.. do you get what im talking about? They're not the same number, 2.31 is rounded off

OpenStudy (anonymous):

yes I do get that that is why I asked because the resulting number would not be the same if you round